### v1

@@ Abstract
Counterfactual image generation is pivotal for understanding the causal relations of variables, with applications in interpretability and generation of unbiased synthetic data. However, evaluating image generation is a long-standing challenge in itself. The need to evaluate counterfactual generation compounds on this challenge, precisely because counterfactuals, by definition, are hypothetical scenarios without observable ground truths. In this paper, we present a novel comprehensive framework aimed at benchmarking counterfactual image generation methods. We incorporate metrics that focus on evaluating diverse aspects of counterfactuals, such as composition, effectiveness, minimality of interventions, and image realism. We assess the performance of three distinct conditional image generation model types, based on the Structural Causal Model paradigm. Our work is accompanied by a user-friendly Python package which allows to further evaluate and benchmark existing and future counterfactual image generation methods [^1] https://github.com/gulnazaki/counterfactual-benchmark. Our framework is extendable to additional SCM and other causal methods, generative models, and datasets.
Keywords:
Benchmark Causal Inference Counterfactual Image Generation
@@ 1 Introduction
“How would I look if I had a beard?", “How a medical scan would be if a patient had disease X?" Counterfactual questions arising from images are very common in science and everyday life. From generating hypothetical visualisations for decision-making to data augmentations for robust model training, counterfactual image generation has emerged as a pivotal domain in the field of artificial intelligence[50, 18].
However, evaluating the efficacy of methods for counterfactual image generation poses a considerable challenge [52]. By definition, counterfactual generation lacks ground truth: e.g., we can’t expect a patient to simultaneously have and not have a disease. This radically complicates the evaluation of counterfactual image generation. Given the importance and frequency with which we move towards counterfactual reasoning in scientific decision-making and everyday life, it is of paramount importance to identify suitable mechanisms for evaluating counterfactual image generation algorithms.
@@
Under a Pearlian methodology of causality, we adopt a set of Structural Causal Models (SCM) [32] to inform generative models about causal relations between factors that affect image generation. Based on the SCM, we employ the counterfactual inference paradigm of Abduction-Action-Prediction and the mechanisms of Deep Structural Causal Models (Deep-SCM) [31] to assess the performance of each method investigated. As such, in this paper, we introduce a comprehensive framework to extensively evaluate any published method on SCM-based counterfactual image generation.
In short, the main contributions of our work are: (1) We develop a comprehensive framework to evaluate the performance of any published image generation method under the Deep-SCM paradigm. While our work focuses on evaluating methods that fall into this category, it can incorporate further/ future causal mechanisms, generative models and datasets. (2) The systematisation of counterfactual image evaluation metrics. (3) We argue in favour of adopting realism and minimality as evaluation metrics for determining successful counterfactuals. (4) The platformisation of the framework in an easy to use python package. (5) The benchmarking and side-by-side comparison of the 3 most prominent families of models used for causal SCM-based counterfactual image generation.
@@ 2 Related work
Causal Counterfactual Image Generation: The pairing of SCM s with deep learning mechanisms for counterfactual image generation can be traced back to the introduction of the Deep-SCM framework [31]. The authors utilise normalising flows [44] and variational inference [20] to infer the exogenous noise and perform causal inference under the no unobserved confounding assumption. Follow-up work [4] incorporates hierarchical latent variable models [2] to improve the fidelity of counterfactuals. In parallel, Generative Adversarial Networks (GANs) [11, 7] have been used to perform counterfactual inference through an adversarial objective [3].
@@
GANs have been previously used in [23] to predict reparametrised distributions over image attributes, but only for interventional inference, while Variational Autoencoders (VAEs) have been used for counterfactual inference [51], focusing on learning structured representations from data, and capturing causal relationships between variables, thus diverging from the above scope. Diffusion models [17, 42] were also recently used for this task [37, 36, 9], approaching the noise abduction as a forward diffusion process. However, the aforementioned work only addresses the scenario of a single variable affecting the image.
@@
Similarly to Deep-SCM, normalising flows and VAEs were used in [22], following a backtracking approach [24]. In backtracking interpretation of counterfactuals, causal mechanisms remain unchanged under intervention, but the exogenous noise is modified. Another line of work, [35], utilises deep twin networks [49] to perform counterfactual inference in the latent space instead of Abduction-Action-Prediction paradigm. Since these methods deviate from our scope and focus on Deep-SCM-based causal counterfactuals, we chose not to include them in the current benchmark. As we aim to explore further causal mechanisms, models and data, we believe that these methods are worthwhile future extensions of our current work.
@@
Evaluation of counterfactuals: To the best of our knowledge, there is no previous work in developing a comprehensive framework to extensively evaluate the performance of counterfactual image generating methods, considering the quality of generated images, as well as their relation to factuals and intervened variables. Our work concentrates on addressing this gap in the literature. A study relevant to ours is [45], which compares established methods for counterfactual explainability [27, 5]. However, the main focus lies on creating counterfactual explanations for classifiers which are done only on one dataset (MorphoMNIST), by solely utilising causal generative models [31, 3].
@@
Although there is no systematic comparison across models and data, various metrics have been proposed for counterfactual image generation. For instance, Monteiro et al. [30] introduce metrics, based on the axiomatic definition of counterfactuals [12, 10], namely the properties that arise from the mathematical formulation of SCMs, which we adopt in our work. The authors of [48] evaluate sparsity of counterfactual explanations via elastic net loss and similarity of factual and counterfactual distributions via autoencoder reconstruction errors. Sanchez & Tsaftaris [37] introduce a metric for evaluating minimality through the latents of a VAE, which we adopt in our work.
Optimising over criteria of causal faithfulness (such as the above), does not directly account for quality of produced images. Evaluation metrics that are used in the broader field of image generation include the Fréchet inception distance (FID) [15] and image similarity metrics such as the Learned Perceptual Image Patch Similarity (LPIPS) [53] or CLIPscore [14]. However, these metrics alone do not suffice for the task of image editing [^2] Image editing has similarities to counterfactual image generation, but does not account necessarily for causal effects between variables. It has gained significant attention recently, with the use of GANs [25, 38] and diffusion models [29, 13] yielding impressive results. We limit our focus to true causal counterfactual methods.; it is common to require spatial information in the form of annotations of expert delineated edit masks information to evaluate a model’s editing abilities [54].
We aspire to explore and include a set of metrics that can consistently evaluate all desired aspects of counterfactual images and serve as proving ground for upcoming methods.
@@ 3.1 Preliminaries
SCM-based counterfactuals The methods we are comparing fall into the common paradigm of SCM-based interventional counterfactuals [19]. A Structural Causal Model (SCM) $\mathcal{G}:=(\mathbf{S},P(\boldsymbol{\epsilon}))$ consists of:
  1. i.
    A collection of structural assignments, called mechanisms $\mathbf{S}=\{f_{i}\}_{i=1}^{N}$, s.t. $x_{i}=f_{i}(\epsilon_{i},\mathbf{pa}_{i})$; and
  2. ii.
    A joint distribution $P(\boldsymbol{\epsilon})=\prod_{i=1}^{N}p(\epsilon_{i})$ over mutually independent noise variables,
where $x_{i}$ is a random variable, $\mathbf{pa}_{i}$ are the parents of $x_{i}$ (its direct causes) and $\epsilon_{i}$ is a random variable (noise). $x_{i}$ and $\mathbf{pa}_{i}$ are observable variables, hence termed endogenous, while $\epsilon_{i}$ is unobservable, exogenous.
@@
Causal relationships are represented by a causal graph in the form of a directed acyclic graph (DAG). Because of this acyclicity, we can recursively solve for $x_{i}$ and obtain a function $\mathbf{x}=\mathbf{f}(\boldsymbol{\epsilon})$. In our context, we assume $\mathbf{x}$ to be a collection of observable variables, where $x_{i}$ can be a high-dimensional object such as an image and $x_{j\neq i}$ attributes of that image.
@@
The assumptions made in the causal models we examine are: (i) The causal graph is known (captured by a SCM). (ii) There exist no unobserved confounders (causal sufficiency), which makes the noise variables $\boldsymbol{\epsilon}$ mutually independent. (iii) The causal mechanisms are invertible, which means that we can write $\epsilon_{i}=f_{i}^{-1}(x_{i},\mathbf{pa}_{i})$ for $i\in[1,n]$ and therefore $\boldsymbol{\epsilon}=\mathbf{f}^{-1}(\mathbf{x})$.

### v5

@@ Abstract
Generative AI has revolutionised visual content editing, empowering users to effortlessly modify images and videos. However, not all edits are equal. To perform realistic edits in domains such as natural image or medical imaging, modifications must respect causal relationships inherent to the data generation process. Such image editing falls into the counterfactual image generation regime. Evaluating counterfactual image generation is substantially complex: not only it lacks observable ground truths, but also requires adherence to causal constraints. Although several counterfactual image generation methods and evaluation metrics exist, a comprehensive comparison within a unified setting is lacking. We present a comparison framework to thoroughly benchmark counterfactual image generation methods. We evaluate the performance of three conditional image generation model families developed within the Structural Causal Model (SCM) framework. We incorporate several metrics that assess diverse aspects of counterfactuals, such as composition, effectiveness, minimality of interventions, and image realism. We integrate all models that have been used for the task at hand and expand them to novel datasets and causal graphs, demonstrating the superiority of Hierarchical VAEs across most datasets and metrics. Our framework is implemented in a user-friendly Python package that can be extended to incorporate additional SCMs, causal methods, generative models, and datasets for the community to build on. Code: https://github.com/gulnazaki/counterfactual-benchmark.
@@ 1 Introduction
Generative AI has revolutionized visual content editing, empowering users to effortlessly modify images and videos (9; 30; 45; 34; 17). However, not all edits are equal, especially in complex domains such as natural or medical imaging, where image realism is paramount (58; 23). In these fields, modifications must adhere to the underlying causal relationships that govern the data generation process. Ignoring these, can lead to unrealistic and potentially misleading results, undermining the integrity of the edited content (46; 58). Image editing inherently relies on implied causal relationships. However, learned generative models powering image editing, will not explicitly account for these causal relations and can produce unrealistic results. Figure 1 showcases clearly what happens when causal relationships are not taken into account and motivates the need for causally faithful approaches.
Image editing, where modifications intentionally alter the underlying causal relationships within the image to explore hypothetical “what if” scenarios, falls squarely into the counterfactual image generation regime (46). This emerging field seeks to generate images that depict plausible yet alternative versions of reality, allowing us to visualize the potential consequences of changes to specific attributes or conditions. Figure 1(c) displays image edits that take into account the causal graph of Figure 1(a), while Figure 1(d) depicts image edits that do not. Intervening on Gender (do(Female)) results in a female person with a beard, while intervening on Age (do(Young)) results in a young person with baldness, which do not account for the causal relationships in the data-generation process.
Evaluating counterfactual image generation is substantially challenging (62; 46). Not only do these hypothetical scenarios lack observable ground truths, but their assessment requires adhering to causal constraints whilst ensuring the generated images are of high quality and visually plausible. This multifaceted problem demands a nuanced approach that balances the need for realism with the constraints imposed by causal relationships.
(a) Causal graph
(b) Factual
(c) Causal
(d) Non-causal
Figure 1: (a) A plausible causal graph for human faces; (b) Factual images (no intervention); (c) Causal counterfactual images using the graph of (a) to perform the interventions $do(Female)$ (upper panel) and $do(Young)$ (lower panel); (d) Non-causal image editing.
@@
Under the Pearlian framework, we adopt a set of Structural Causal Models (SCM) (38) to explicitly inform generative models about causal paths between high-level variables. Based on the SCM, we employ the counterfactual inference paradigm of Abduction-Action-Prediction and the mechanisms of Deep Structural Causal Models (Deep-SCM) (36) to evaluate the performance of various methods. As such, in this paper, we introduce a comprehensive framework, able to extensively evaluate any published methods on SCM-based counterfactual image generation. We aspire for this work to become the de facto benchmark for the community when developing new methods (and metrics) in counterfactual image generation.
Briefly, the main contributions of our work are: (1) We develop a comprehensive framework to evaluate the performance of all image generation models published under the Deep-SCM paradigm. We explore several datasets (synthetic, natural and medical images), as well as SCM structures across datasets. (2) We expand the above models to accommodate previously untested datasets and causal graphs. Specifically, we test HVAE and GAN on a non-trivial causal graph for human faces and we devise a GAN architecture inspired from (59) that can generate counterfactual brain MRIs given multiple variables. (3) We extensively benchmark these models adopting several metrics to evaluate causal SCM-based counterfactual image generation. (4) We offer a user-friendly Python package to accommodate and evaluate forthcoming causal mechanisms, datasets and causal graphs.
@@ 2 Related work
Causal Counterfactual Image Generation: Pairing SCM s with deep learning mechanisms for counterfactual image generation can be traced back to the emergence of Deep-SCM (36). The authors utilised Normalising Flows (51) and variational inference (25) to infer the exogenous noise and perform causal inference under the no unobserved confounders assumption. Follow-up work (7) incorporates a Hierarchical VAE (HVAE) (5) to improve the fidelity of counterfactuals. In parallel, Generative Adversarial Networks (GANs) (14; 10) have been used to perform counterfactual inference through an adversarial objective (6) and for interventional inference, by predicting reparametrised distributions over image attributes (28). Variational Autoencoders (VAEs) have been used for counterfactual inference (61), focusing on learning structured representations from data and capturing causal relationships between high-level variables, thus diverging from the above scope. Diffusion models (21; 48), on the other hand, were also recently used to learn causal paths between high-level variables (44; 43; 12), approaching the noise abduction as a forward diffusion process. However, these works (44; 43; 12) address the scenario of a single variable (attribute) affecting the image.
@@
Similarly to Deep-SCM, Normalising Flows and VAEs were used in (27), following a backtracking approach (29). The aforementioned method involves tracing back through the causal graph to identify changes that would lead to the desired change in the attribute intervened. Another line of work, (42), utilises deep twin networks (57) to perform counterfactual inference in the latent space, without following the Abduction-Action-Prediction paradigm. Since these methods deviate from our scope and focus on Deep-SCM-based causal counterfactuals, we chose not to include them in the current benchmark. As we aim to explore further causal mechanisms, models and data, we believe that these methods are worthwhile future extensions of our current work.
Recent works based on diffusion models (21) have achieved SoTA results on image editing (22; 34; 17; 2; 60; 54), usually by altering a text prompt to edit a given image. While the produced images are of high quality, we aim to explicitly model the causal interplay of known high-level variables, an approach not yet applied to diffusion models to the best of our knowledge. Therefore, even though the task of image editing exhibits similarities to counterfactual image generation, we focus on true causal counterfactual methods, as we discussed above.
@@
Evaluation of counterfactuals: To the best of our knowledge, there is no prior work offering a comprehensive framework to thoroughly evaluate the performance, fidelity, and reliability of counterfactual image generation methods, considering both image quality and their relationship to factual images and intervened variables. This paper aims to address this gap. A study relevant to ours is (52): it compares established methods for counterfactual explainability [^1] Counterfactual explanations aim to explain changes in the predictions of classifiers through minimal changes of the input. Although both our task and counterfactual explainability can benefit from each other, they differ in terms of their objective and methodology. (8; 32). However, its main focus lies on creating counterfactual explanations for classifiers on a single dataset (MorphoMNIST (4)) using solely two causal generative models (VAE, GAN) (36; 6).
@@
Various metrics have been proposed to evaluate counterfactual image generation. For instance, Monteiro et al. (35) introduce metrics, following an axiomatic definition of counterfactuals (15; 13), namely the properties that arise from the mathematical formulation of SCMs. The authors in (56) evaluate sparsity of counterfactual explanations via elastic net loss and similarity of factual versus counterfactual distributions via autoencoder reconstruction errors. Sanchez and Tsaftaris (44) introduce a metric to evaluate minimality in the latent space. In this work, we adopt the metrics of (35) and (44).
Optimising criteria of causal faithfulness as those above does not directly ascertain image quality. Image quality evaluation metrics used broadly in the context of image editing include the Fréchet inception distance (FID), (19) and similarity metrics such as the Learned Perceptual Image Patch Similarity (LPIPS) (63) and CLIPscore (18). We adopt such image quality evaluation metrics herein.
@@ 3.1 Preliminaries
SCM-based counterfactuals The methods we are comparing fall into the common paradigm of SCM-based interventional counterfactuals (37). A Structural Causal Model (SCM) $\mathcal{G}:=(\mathbf{S},P(\boldsymbol{\epsilon}))$ consists of: (i) A collection of structural assignments, called mechanisms $\mathbf{S}=\{f_{i}\}_{i=1}^{N}$, s.t. $x_{i}=f_{i}(\epsilon_{i},\mathbf{pa}_{i})$; and (ii) A joint distribution $P(\boldsymbol{\epsilon})=\prod_{i=1}^{N}p(\epsilon_{i})$ over mutually independent noise variables, where $x_{i}$ is a random variable, $\mathbf{pa}_{i}$ are the parents of $x_{i}$ (its direct causes) and $\epsilon_{i}$ is a random variable (noise).
@@
Causal relations are represented by a directed acyclic graph (DAG). Due to the acyclicity, we can recursively solve for $x_{i}$ and obtain a function $\mathbf{x}=\mathbf{f}(\boldsymbol{\epsilon})$. We assume $\mathbf{x}$ to be a collection of observable variables, where $x_{i}$ can be a high-dimensional object such as an image and $x_{j\neq i}$ image attributes. The variables $\mathbf{x}$ are observable, hence termed endogenous, while $\boldsymbol{\epsilon}$ are unobservable, exogenous.
Figure 2: Inference example on MorphoMNIST: Abduction: We abduct the exogenous noise separately for each variable, using inverse Normalising Flows $f^{-1}$ for the attributes and the encoder of the image mechanism (e.g. VAE, HVAE or GAN) for the image, conditioning on the factual parents. Action: We intervene ($do(t^{*})$) only on thickness. Prediction: We employ the Normalizing Flow $f$ conditioned on the counterfactual thickness $t^{*}$ to obtain $i^{*}$ after the intervention. Note that this is not needed for $t^{*}$ on which we intervene and for $d^{*}$ that has no parents in the DAG. Finally, the decoder generates the counterfactual image, given the exogenous noise $U_{img}$ and conditioning on the counterfactual parents.
@@
The models we examine make several assumptions: (i) Known causal graph (captured by a SCM). (ii) Lack of unobserved confounders (causal sufficiency), making the $\epsilon_{i}$, $\epsilon_{j}$ mutually independent. (iii) Invertible causal mechanisms, i.e. $\epsilon_{i}=f_{i}^{-1}(x_{i},\mathbf{pa}_{i})$ $\forall$ $i\in[1,n]$ and hence $\boldsymbol{\epsilon}=\mathbf{f}^{-1}(\mathbf{x})$.